OpenStudy (anonymous):

The three sides of a right triangle form a geometric sequence. What is the ratio of the length of the shortest side to the length of the hypotenuse? Express your answer as a decimal to the nearest tenth.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (jdoe0001):

btw, is there a picture you'd want to show us?

OpenStudy (anonymous):

No, there is no pic :(

OpenStudy (anonymous):

Sorry, I am not sure about the link you sent...

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

Now I kinda understand the link, jdoe

OpenStudy (anonymous):

problem is, I don't have any measurements

OpenStudy (jdoe0001):

the exercise assumes there is a "shortest side" and a ratio to the hypotenuse, and expects a rounded answer, thus there's some right angle we're not seeing

OpenStudy (anonymous):

um...

OpenStudy (anonymous):

does anyone have an idea how to solve or does anyone think it's imposs?

OpenStudy (cwrw238):

lets say the three lengths are a , ar and ar^2 then required ratio is a / ar^2 = 1 / r^2 also a^2r^4 = a^2 + a^2r^2 r^4 = r^2 + 1 let R = r^2 R^2 - R - 1 = 0 R = r^2 = 1.618 ratio = 1 /1.618 i think thats right

OpenStudy (anonymous):

Wow! You actually did that!

OpenStudy (anonymous):

i think i understand now...thx

OpenStudy (anonymous):

i'll look it over but it looks reasonable to me:)

OpenStudy (cwrw238):

- I used formula for terms of a GP and also Pythagoras this only applies when the right angles triangle has different lengthed sides

OpenStudy (cwrw238):

yea i think it makes sense

OpenStudy (anonymous):

I noticed the Pythagoras

OpenStudy (anonymous):

what's gp stand for?

OpenStudy (anonymous):

GP

OpenStudy (cwrw238):

suppose we take a to be 2 and work out ar and ar^2 and see it they fir GP = gemetric progression ( or series)

OpenStudy (anonymous):

(or sequence) yep

OpenStudy (cwrw238):

yes

OpenStudy (cwrw238):

if a = 2 the 3 sides will have length 2, 2*sqrt1.618 and 2*1.618 by pythagoras (2*1.618)^2 = 2^2 + (2*sqrt1.618)^2 10.472 = 10.472 it checks out